OpenStudy (anonymous):

Natalie picked 135 berries in 15 minutes. If she continues picking at that rate, how long will it take her to pick 486 berries?

hero (hero):

\[\frac{135 \text{berries}}{486 \text{berries}} = \frac{15 \min}{x \min}\]

hero (hero):

\[\frac{135 \cancel{\text{berries}}}{486 \cancel{\text{berries}}} = \frac{15 \cancel{\min}}{x \cancel{\min}}\]

hero (hero):

\[\frac{135}{486}= \frac{15}{x}\]

hero (hero):

Cross Multiply

OpenStudy (anonymous):

ok thankyou